OpenStudy (anonymous):

What is the equation of a vertical line passing through (–4, –3)?

OpenStudy (anonymous):

x = –4 y = –4 x = –3 y = –3

OpenStudy (anonymous):

@ivettef365

OpenStudy (ivettef365):

value for x = vertical value for y = horizontal

OpenStudy (ivettef365):

so it is vertical and what is the value of x

OpenStudy (ivettef365):

you have (x,y) point (-4,-3)

OpenStudy (anonymous):

ok :) i got to that part, but which one do i pick?

OpenStudy (anonymous):

@ivettef365

OpenStudy (ivettef365):

what is the x value on point (-4,-3) ??

OpenStudy (anonymous):

-4

OpenStudy (ivettef365):

ok so x = -4

OpenStudy (anonymous):

Simplify (x – 4)(x2 – 2x – 3)

OpenStudy (anonymous):

how about this one?

OpenStudy (anonymous):

@ivettef365

OpenStudy (ivettef365):

you need to multiply as follows: x(x^2-2x-3)-4(x^2-2x-3)

OpenStudy (anonymous):

i only have like 5 more questions. can you help witht them?

OpenStudy (ivettef365):

sure

OpenStudy (ivettef365):

can you solve the above equation

OpenStudy (anonymous):

x3-6x2-3x - 4x2-8x-12?

OpenStudy (anonymous):

x3 – 6x2 + 5x - 12

OpenStudy (anonymous):

?

OpenStudy (anonymous):

multiply (2sqrt3+3)(3sqrt3-5) <---- another one

OpenStudy (anonymous):

x3 – 6x2 + 5x + 12 or this?

OpenStudy (ivettef365):

x^3 -2x^2 -3x -4x^2 +8x +12

OpenStudy (anonymous):

x3 – 6x2 + 5x + 12 so this one?

OpenStudy (ivettef365):

correct

OpenStudy (anonymous):

you should be the one getting all the metals. you always help

OpenStudy (anonymous):

multiply (2sqrt3+3)(3sqrt3-5) @ivettef365

OpenStudy (ivettef365):

ok 2\(\sqrt{3}\) . 3\(\sqrt{3}\) = 6.3 = 18 2\(\sqrt{3}\) . (-5) = -10\(\sqrt{3}\) (3) . 3\(\sqrt{3}\) = 9\(\sqrt{3}\) (3) . (-5) = -15

OpenStudy (ivettef365):

so you have 18 - 10\(\sqrt{3}\) + 9\(\sqrt{3}\) -15

OpenStudy (ivettef365):

can you take it from there

OpenStudy (anonymous):

:/

OpenStudy (ivettef365):

just combine like terms

OpenStudy (anonymous):

3- sqrt3

OpenStudy (ivettef365):

correct

OpenStudy (anonymous):

Choose the system of equations which matches the following graph.

OpenStudy (anonymous):

x + 2y = 4 3x – 6y = 18 x – 2y = 4 3x + 6y = 18 x + 2y = 4 3x + 6y = 18 x – 2y = 4 3x – 6y = 18

OpenStudy (anonymous):

wont they be negative?

OpenStudy (ivettef365):

it can not be a, because when you convert to y = mx + b, where m = slope you can see that both lines have negative slope and are parallel, so the slope needs to be negative and the same. you have on a) x + 2y = 4 2y = -x + 4 y = -1/2x +2 and 3x - 6y = 18 -6y = -3x + 18 (divide by -6) you have y = 1/2 x - 3 one have slope of -1/2 the other slope of 1/2 not good

OpenStudy (anonymous):

d then

OpenStudy (ivettef365):

no, lets try c x + 2y = 4 2y = -x + 4 y = -1/2x + 2 (y-int = 2) negative slope is good 3x + 6y = 18 6y = -3x + 18 y = -1/2x + 3 (negative slope, same slope, y-int = 3) is good

OpenStudy (anonymous):

oh ok :) im so dumb

OpenStudy (ivettef365):

no, you're not, is just a matter of practice

OpenStudy (anonymous):

What is the slope of the line which passes through (–2, 0) and (0, 4)? how about this

OpenStudy (anonymous):

–2 0 2 Undefined

OpenStudy (anonymous):

is it d

OpenStudy (ivettef365):

that is just using the slope formula y2 - y1 ------ x2 - x1

OpenStudy (anonymous):

its undefined

OpenStudy (anonymous):

it 2!

OpenStudy (ivettef365):

try this website, enter the points and see what you get

OpenStudy (anonymous):

4/2

OpenStudy (ivettef365):

correct, which is 2

OpenStudy (anonymous):

a couple more. multiply x^5/6 x^3/4

OpenStudy (anonymous):

@ivettef365

OpenStudy (anonymous):

@hobbs978

OpenStudy (ivettef365):

when multiplying exponents, you need to add so you have x^5/6 . x^3/4 so it is equal to x^(5/6 + 3/4)

OpenStudy (anonymous):

x4/5

OpenStudy (anonymous):

Simplify the expression

OpenStudy (anonymous):

@ivettef365

OpenStudy (ivettef365):

5 3 5(2) + 3(3) 19 --- + ---- = ---------- = -------- 6 4 12 12 SO IT IS X^19/12

OpenStudy (anonymous):

not an option

OpenStudy (ivettef365):

are you sure you posted the right problem, you have multiply x^5/6 . x^3/4

OpenStudy (ivettef365):

the next one you have \(\sqrt{-9}\) that is equal to 3i

OpenStudy (ivettef365):

so you have 3i ------------- 3 - 2i + 1 + 5i (combine like terms, what do you get)

OpenStudy (anonymous):

4+3i

OpenStudy (ivettef365):

so the answer is 3i --------- 4 + 3i

OpenStudy (ivettef365):

did you find out about the other problem ??

OpenStudy (anonymous):

Simplify i37

OpenStudy (anonymous):

i think its i